# Does quantum information require additional structure?

### v2

@@ 8.1 Internal relation quantum space time $\mathbf{Q}_{rel,4}$
Now consider the directly measurable observables $\widetilde{G}(p_{\mu}^{2})=\widetilde{G^{\prime}}(p_{\mu}^{\prime 2})$, assuming that their $x$ representations are well defined, i.e., the corresponding integrals converge, in the asymptotic zone $M$-form-invariant function. Then the $M_{4}-x$ representation of Hilbert vector $|G\rangle$ is given by
@@
where $\xi_{\mu}$ are two arbitrary four points (events) in $M_{4}$ and $p_{\mu}$ are relative 4-momenta of the whole fully isolated system. Here we assume that (13) is well determined under a proper contour $D$ in the $p_{0}$ – complex plane.
Definition 1.
The three-dimensional Lorentz absolute momentum space $\mathbf{Q}_{rel,3}$ spanned by the momentum $\mathbf{q}$ is defined as a space in which the auxiliary function $\tilde{F}(q^{2})$ is embedded so that, for spacelike $p_{\mu}$
$\widetilde{F}(\mathbf{q}^{2})=\widetilde{G}(p_{\mu}^{2})=p_{\mu}^{\prime 2}=\mathbf{q}^{2}\geq 0)$ (12)
Here $p_{\mu}^{2}$, $p_{\mu}^{\prime 2}$ are representations of the 4-momentum in arbitrary frames $S,S$. Note that the above equality which defines internal space $\mathbf{Q}_{rel,3}$ does not distinguish any reference frame $S$ in $M_{4}$. By definition $\widetilde{F}(\mathbf{q}^{2})$ is implicitly Lorentz-form-invariant whereas $\widetilde{G}(p_{\mu}^{2})$ is explicitly Lorentz-form-invariant. In this way, the directly measurable functions $G$ implicitly the determine $L$ absolute 3-dimensional space $Q_{rel,3}$. In accordance with the quantum canonical symmetry $\tilde{F}(q)$ is interpreted as the momentum representation in $\mathbf{Q}_{rel}$ of the Hilbert vector $F$ space $|F\rangle$, $\widetilde{F}(\mathbf{q}^{2})=\langle\mathbf{q}|F\rangle$
@@
Note that the class of $|F\rangle$’s is restricted according to the convergence of the corresponding integrals. In particular interval $|\mathbf{y}|=F(\mathbf{y}^{2})=r$ in $\mathbf{Q}_{rel,3}$ has no representation in $M_{4}$ while the interval $x^{2}=G(x^{2})$ has no representation in $\mathbf{Q}_{rel,3}$. However, all typical bound states and Yukawa-type potentials have both internal and external representations in $\mathbf{Q}_{rel,4}$ and $L_{4}$.
Now, according to the quantum canonical symmetry $\mathbf{Q}$ in the Schrödinger representation $\hat{\mathbf{y}}=\mathbf{y}$, $\hat{\mathbf{q}}=-i\hbar\textnormal{grad}_{y}$ one can introduce an $M$-absolute internal Hamiltonian operator $\widehat{h}(\hat{\mathbf{y}},\hat{\mathbf{q}})$ in the configuration space $\mathbf{Q}_{rel,3}$, and the corresponding Schrödinger equation of the internal motion of an isolated system $S_{M}$ can be written in the form:
@@
where the internal Hamiltonian $\hat{h}$ is spanned by $\hat{\mathbf{y}}$ and $\hat{\mathbf{q}}$ and, consistent with the relativistic kinematics, defines the absolute evolution parameter $\tau$ which completes the space $\mathbf{Q}_{rel,4}$ into the 4-dimensional continuum of the internal space-time $S_{M}$ and hence it is the Cartesian product $\mathbf{Q}_{rel,4}=\mathbf{Q}_{rel,3}\times T$ which operates with the two absolute intervals $r=|\mathbf{y}|$, $\Delta\tau=\Delta t$, which coincide with the geometry intervals $\mathbf{Q}_{rel,4}$ (10). Thus the nonrelativistic limit of $\mathbf{Q}_{rel,3}$ is the space $\mathbf{Q}_{rel,3}^{G}$ immersed in the Galilean space-time, while the internal time $\tau$ coincides up to an arbitrary translation constant with the $G$-absolute Newtonian time.
@@
It is interesting that this very rough model, ignoring the quark structure of the nucleon for elastic form factor electron–nucleon collisions model predicts proton mean-square radius 0.81fm, while recent experiments provide values: 0.843fm [106], 0.833fm [107, 108] 0.8414fm [109]. These experimental discrepancies indicate that the proton radius problem still remains open.
@@ 9 Conclusion
The concept of quantum information is central to the quantum model. In particular, it is shown that “many fundamental features of quantum systems can be understood as consequences of the constraints that limited information content imposes on the systems” [46]. However, this does not translate easily into the interpretive problems of quantum theory. We try to determine the status of quantum information and its place based on two concepts. The first is the correspondence principle between the physical structure represented by the set of observer-independent data, including fundamental constants of nature, and the mathematical structure of the model that it describes. This principle allows us to interpret the wave function as a mathematical representation (image) of quantum information that is relational in nature. Consequently, we introduced the concept of relational space, in which quantum relations precede correlations, and we postulated that its mathematical structure is isomorphic to the tensorial Hilbert space structure. This interpretation becomes particularly interesting in the context of the inconsistency between Born’s interpretation of the wave function in lab space and von Neumann’s configuration space. In this approach, correlations are interpreted as limiting relations requiring decoherence and irreversibility of recorded results. Hence, this approach naturally excludes Wigner’s type paradoxes. Finally, we consider Chyliński’s quantum relational continuum space model, which leads to potentially observable effects in elastic, inelastic, and dissociation processes only for bound states.
Unfortunately, Chyliński’s articles went unnoticed, and to our knowledge, the theoretical predictions have not been tested experimentally. Even though 40 years have passed, the idea of Chyliński’s internal quantum space still seems interesting, especially in the context of the quantum correlations phenomena.
@@
Acknowledgments The author would like to especially thank L. Newelski for his suggestion to replace the isomorphic principle with a correspondence principle that goes beyond a purely mathematical context. Special thanks are due to A. Wójcik for a panoramic, deep insight into the idea of this paper, and M. Eckstein and K. Horodecki for their deep and constructive criticism. The author would also like to thank M. Eckstein and P. Horodecki for a sequence of inspiring discussions and also M. Czachor and M. Nowakowski for their discussion, and Ł. Pankowski for help with editing the manuscript. The work was partly supported by the Foundation for Polish Science (IRAP Project, ICTQT, Contract No. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme).

### v3

@@ 7.1 Internal Relation Quantum Space Time $\mathbf{Q}_{rel,4}$
Now consider the directly measurable observables $\widetilde{G}(p_{\mu}^{2})=\widetilde{G^{\prime}}(p_{\mu}^{\prime 2})$, assuming that their $x$ representations are well defined, i.e., the corresponding integrals converge, in the asymptotic zone $L$-form-invariant function. Then the $M_{4}-x$ representation of Hilbert vector $|G\rangle$ is given by
@@
where $\xi_{\mu}$ are two arbitrary four points (events) in $M_{4}$ and $p_{\mu}$ are relative 4-momenta of the whole fully isolated system. Here we assume that Eq. 11 is well determined under a proper contour $D$ in the $p_{0}$ – complex plane.
Definition 1
[39] The three-dimensional Lorentz absolute momentum space $\mathbf{Q}_{rel,3}$ spanned by the momentum $\mathbf{q}$ is defined as a space in which the auxiliary function $\tilde{F}(q^{2})$ is embedded so that, for spacelike $p_{\mu}$
$\widetilde{F}(\mathbf{q}^{2})=\widetilde{G}(p_{\mu}^{2})=p_{\mu}^{\prime 2}=\mathbf{q}^{2}\geq 0$ (12)
Here $p_{\mu}^{2}$, $p_{\mu}^{\prime 2}$ are representations of the 4-momentum in arbitrary frames $S,S^{\prime}$. Note that the above equality which defines internal space $\mathbf{Q}_{rel,3}$ does not distinguish any reference frame $S$ in $M_{4}$. By definition $\widetilde{F}(\mathbf{q}^{2})$ is implicitly Lorentz-form-invariant whereas $\widetilde{G}(p_{\mu}^{2})$ is explicitly Lorentz-form-invariant. In this way, the directly measurable functions $G$ implicitly the determine $L$ absolute 3-dimensional space $Q_{rel,3}$. In accordance with the quantum canonical symmetry $\tilde{F}(q)$ is interpreted as the momentum representation in $\mathbf{Q}_{rel}$ of the Hilbert vector $F$ space $|F\rangle$, $\widetilde{F}(\mathbf{q}^{2})=\langle\mathbf{q}|F\rangle$
@@
Note that the class of $|F\rangle$’s is restricted according to the convergence of the corresponding integrals. In particular interval $|\mathbf{y}|=F(\mathbf{y}^{2})=r$ in $\mathbf{Q}_{rel,3}$ has no representation in $M_{4}$ while the interval $x^{2}=G(x^{2})$ in $M_{4}$ has no representation in $\mathbf{Q}_{rel,4}$. However, all typical bound states and Yukawa-type potentials have both internal and external representations in $\mathbf{Q}_{rel,4}$ and $M_{4}$.
Now, according to the quantum canonical symmetry $\mathbf{Q}$ in the Schrödinger representation $\hat{\mathbf{y}}=\mathbf{y}$, $\hat{\mathbf{q}}=-i\hbar\overline{\nabla}_{y}$ one can introduce an $L$-absolute internal Hamiltonian operator $\widehat{h}(\hat{\mathbf{y}},\hat{\mathbf{q}})$ in the configuration-space $\mathbf{Q}_{rel,3}$, and the corresponding Schrödinger equation of the internal motion of an isolated system $S_{M}$ can be written in the form:
@@
where the internal Hamiltonian $\hat{h}$ is spanned by $\hat{\mathbf{y}}$ and $\hat{\mathbf{q}}$ and, consistent with the relativistic kinematics, defines the absolute evolution parameter $\tau$ which completes the space $\mathbf{Q}_{rel,4}$ into the 4-dimensional continuum of the internal space-time $S_{M}$ and hence it is the Cartesian product $\mathbf{Q}_{rel,4}=\mathbf{Q}_{rel,3}\times T$ which operates with the two absolute intervals $r=|\mathbf{y}|$, $\Delta\tau=\Delta t$, which coincide with the geometry intervals $\mathbf{Q}_{rel,4}$ Eq. 10. Thus the nonrelativistic limit of $\mathbf{Q}_{rel,3}$ is the space $\mathbf{Q}_{rel,3}^{G}$ immersed in the Galilean space-time, while the internal time $\tau$ coincides up to an arbitrary translation constant with the $G$-absolute Newtonian time.
@@
It is interesting that this very rough model, ignoring the quark structure of the nucleon for elastic form factor electron–nucleon collisions model predicts proton mean-square radius 0.81fm [111], while recent experiments provide values: 0.843fm [112], 0.833fm [113, 114] 0.8414fm [115]. These experimental discrepancies indicate that the proton radius problem still remains open.
@@ 8 Conclusion
The concept of quantum information is central to the quantum model. In particular, it is shown that “many fundamental features of quantum systems can be understood as consequences of the constraints that limited information content imposes on the systems” [42]. However, this does not translate easily into the interpretational problems of quantum theory. We tried to determine the status of quantum information and its place in quantum theory within the QRS hypothesis. In this aim we have introduced the correspondence principle between the physical structure represented by the set of observer-independence data, and the mathematical structure of the model that it describes. This principle allows us to interpret the wave function as a mathematical representation (image) of quantum information that is relational in nature. Consequently we have introduced an internal $Q_{rel}$ space which precedes the space of correlation events $Q_{cor}$. We assumed that the mathematical structure of this hypothetical space is isomorphic to the tensorial Hilbert space structure and pointed out limitations of this assumption. In this approach quantum correlations are interpreted as limiting relations requiring decoherence and irreversibility of recorded results. Hence, the QRS hypothesis naturally excludes Wigner’s type paradoxes. We discussed some possible consequences of QRS hypothesis, in particular, we showed that assumption that quantum relations precede quantum correlations is compatible with recent experimental and theoretical results [50, 51].
Finally, we consider Chyliński’s quantum relational continuum space model, which leads to potentially observable effects in elastic, inelastic, and dissociation processes only for bound states. Unfortunately, Chyliński’s articles went unnoticed, and to our knowledge, the theoretical predictions have not been tested experimentally. Even though 40 years have passed, the idea of Chyliński’s internal quantum space still seems interesting, especially in the context of the quantum correlations phenomena. In particular, it would be interesting to extend Chyliński’s model to include the spin degrees of freedom.
We believe that our approach, when properly developed, will lead to a deeper understanding of the role of quantum information in quantum theory.
@@
Acknowledgments The author would like to especially thank L. Newelski for his suggestion to replace the isomorphic principle with a correspondence principle that goes beyond a purely mathematical context. Special thanks are due to A. Wójcik for a panoramic, deep insight into the idea of this paper, and M. Eckstein and K. Horodecki for their deep and constructive criticism. The author would also like to thank M. Eckstein and P. Horodecki for a sequence of inspiring discussions and also M. Czachor, A. Grudka, A. Miranowicz and M. Nowakowski for their discussion, and Ł. Pankowski for help with editing the manuscript. Finally, I would like to thank the anonymous Reviewers for their searching and inspiring criticism.
[…]
Funding The work was partly supported by the Foundation for Polish Science (IRAP Project, ICTQT, Contract No. MAB/2018/5, co-financed by the EU within the Smart Growth Operational Programme).